Consider an electricity market where there are three suppliers, each with constant marginal cost (a reasonable approximation in electricity generation). Firm 1 has a capacity of 200 at MC = 5. Firm 2 has a capacity of 100 at MC = 8. Firm 3 has a capacity of 100 at MC = 10.

  1. Ifthethreefirmsarepricetakers(i.e.,behavecompetitively),whatistheindustrysupplycurve?
  2. ComputetheequilibriumpriceandquantitywhenthemarketdemandisQ(P)=1000/P0.5.
  3. Compute the equilibrium price and quantity when the market demand is Q(P) = 750/P0.5.
  4. ComputetheequilibriumpriceandquantitywhenthemarketdemandisQ(P)=500/P0.5.
1
Expert's answer
2021-03-18T00:19:49-0400

1) For the first firm we have the supply curve "Q_{s1}=40P," for the second firm - "Q_{s2}=12.5P" and for the third one - "Q_{s3}=10P." Then, we can find the industry supply curve:


"Q_s=Q_{s1}+Q_{s2}+Q_{s3},""Q_s=40P+12.5P+10P=52.5P."

2) Market equilibrium occurs when "Q_p=Q_s":


"\dfrac{1000}{\sqrt{P}}=52.5P,""\dfrac{10^6}{P}=2756.25P^2,""P=\sqrt[3]{\dfrac{10^6}{2756.25}}=\$7.13.""Q=\dfrac{1000}{\sqrt{\$7.13}}=374.5."

3)

"\dfrac{750}{\sqrt{P}}=52.5P,""\dfrac{562500}{P}=2756.25P^2,""P=\sqrt[3]{\dfrac{562500}{2756.25}}=\$5.89.""Q=\dfrac{750}{\sqrt{\$5.89}}=309."

4)

"\dfrac{500}{\sqrt{P}}=52.5P,""\dfrac{250000}{P}=2756.25P^2,""P=\sqrt[3]{\dfrac{250000}{2756.25}}=\$4.49.""Q=\dfrac{500}{\sqrt{\$4.49}}=236."